# Learning Theory for Kernel Bilevel Optimization

### v1

@@ Abstract
Bilevel optimization has emerged as a technique for addressing a wide range of machine learning problems that involve an outer objective implicitly determined by the minimizer of an inner problem. In this paper, we investigate the generalization properties for kernel bilevel optimization problems where the inner objective is optimized over a Reproducing Kernel Hilbert Space. This setting enables rich function approximation while providing a foundation for rigorous theoretical analysis. In this context, we establish novel generalization error bounds for the bilevel problem under finite-sample approximation. Our approach adopts a functional perspective, inspired by (Petrulionytė et al., 2024), and leverages tools from empirical process theory and maximal inequalities for degenerate $U$-processes to derive uniform error bounds. These generalization error estimates allow to characterize the statistical accuracy of gradient-based methods applied to the empirical discretization of the bilevel problem.
@@ 1 Introduction
Optimization is a cornerstone of mathematical modeling, machine learning, and decision-making. Among the various frameworks, bilevel optimization stands out due to its hierarchical structure, where one optimization problem, called outer-level, is constrained by the solution of another problem, called inner-level (Candler & Norton, 1977). This framework was first introduced in the context of economic game theory by von Stackelberg (1934) to model leader-follower interactions, where one agent’s decisions depend on the optimal response of another. Over time, bilevel optimization has naturally found applications in a broad spectrum of machine learning fields, including hyper-parameter tuning (Larsen et al., 1996; Bengio, 2000; Franceschi et al., 2018), meta-learning (Bertinetto et al., 2019; Pham et al., 2021), inverse problems (Holler et al., 2018), and reinforcement learning (Hong et al., 2023; Liu et al., 2023), making it a powerful and versatile tool in both theoretical and practical contexts.
@@
The success of bilevel optimization in machine learning naturally raises fundamental questions about the generalization properties of models learned through these procedures, as the number of data samples increases. Several existing works have studied the generalization and convergence of bilevel algorithms under the assumption that the inner-level problem is strongly convex and that its parameters lie in a finite-dimensional space. These include analyses of the convergence of stochastic bilevel optimization algorithms (Arbel & Mairal, 2022a; Dagréou et al., 2022; Ghadimi & Wang, 2018; Ji et al., 2021) and approaches based on algorithmic stability (Bao et al., 2021; Zhang et al., 2024). The strong convexity assumption ensures the uniqueness of the inner-level solution, a key property for stability and convergence analysis in bilevel optimization. Moreover, restricting the inner-level parameters to a finite-dimensional space, instead of possibly richer infinite-dimensional spaces, as in kernel methods, circumvents additional complexities, where the parameter’s dimension may grow with the sample size. This dependence on sample size, in non-parametric methods, poses additional challenges, as solutions at different sample sizes are not directly comparable. In contrast, in the finite-dimensional setting, generalization bounds can be derived by quantifying the convergence of finite-sample estimates of the inner-level solution toward the population solution, i.e., the solution obtained in the limit of infinite samples, within the same parameter space.
@@
Albeit convenient from a theoretical perspective, having both strong convexity and finite-dimensionality drastically limits expressiveness of the models, effectively restricting them to linear functions. Going beyond linear models requires either relaxing the strong convexity assumption to accommodate more expressive models, such as deep neural networks (Goodfellow et al., 2016), or considering non-parametric bilevel problems, where the inner-level variable lies in an expressive infinite-dimensional function space, such as a Reproducing Kernel Hilbert Space (RKHS) (Schölkopf & Smola, 2002). Early works in bilevel optimization for machine learning followed the latter approach, developing bilevel methods for hyper-parameter selection in the context of kernel methods (Keerthi et al., 2006; Kunapuli et al., 2008). These works leverage the representer theorem (Schölkopf et al., 2001) to transform the infinite-dimensional problem into a finite-dimensional one with dimension depending on the sample size. However, they do not address how the sample size impacts generalization. Another line of research instead focuses on relaxing the strong convexity assumption, proposing new bilevel algorithms that can handle the loss of convexity (Arbel & Mairal, 2022b; Kwon et al., 2024; Shen & Chen, 2023). Nevertheless, non-convex bilevel optimization is a very hard problem in general (Liu et al., 2021; Arbel & Mairal, 2022b; Bolte et al., 2024), and obtaining strong generalization guarantees in this setting remains out of reach due to the lack of precise control over the inner-level solution. In all cases, learning theory for bilevel problems beyond the strongly convex parametric setting is essentially lacking.
@@
In the present work, we take an initial step towards developing a learning theory that goes beyond the finite-dimensional setting. Specifically, we propose to study Kernel Bilevel Optimization (KBO) problems, where the inner objective $L_{in}:\mathbb{R}^{d}\times\mathcal{H}\rightarrow\mathbb{R}$ finds an optimal inner solution $h_{\omega}^{\star}$ in a RKHS $\mathcal{H}$ for a given parameter $\omega$ in $\mathbb{R}^{d}$, while the outer objective $L_{out}:\mathbb{R}^{d}\times\mathcal{H}\rightarrow\mathbb{R}$ optimizes the parameter $\omega$ over a closed subset $\mathcal{C}$ of $\mathbb{R}^{d}$, given the inner solution $h_{\omega}^{\star}$:
$\min_{\omega\in\mathcal{C}}\mathcal{F}(\omega)\coloneqq L_{out}(\omega,h^{\star}_{\omega})$ (KBO)
[…]
@@
In particular, we focus on inner and outer level objectives that are expectation of point-wise losses, a common setting in learning theory. RKHS provides a natural framework to study learning theoretic arguments, and has been instrumental for many fruitful results in pattern recognition and machine learning. They allow to describe very expressive non-linear models with simple and stable algorithms, while enabling a rich statistical analysis and featuring adaptivity to the regularity of the population problem (Shawe-Taylor & Cristianini, 2004; Schölkopf & Smola, 2002; Hofmann et al., 2008). Our choice is also motivated by the relevance of kernel methods, even in the deep learning era. They remain competitive for some prediction problem, for example involving physics (Doumèche et al., 2024; Letizia et al., 2022). Furthermore, the mathematics of kernel methods are useful to describe the limiting behavior of deep network training for very large models (Jacot et al., 2018; Belkin et al., 2018). Indeed, in this limit, the problem becomes (strongly) convex in an infinite-dimensional functional space, simplifying the difficulties of non-convex model parameterization, a major bottleneck in the analysis of such models. This point of view was leveraged by Petrulionytė et al. (2024) who introduced functional bilevel optimization, and our setting is a special case for which the underlying function space is an RKHS. From a practical perspective, our setting is amenable to first-order methods using implicit differentiation techniques (Griewank & Faure, 2003; Bai et al., 2019; Blondel et al., 2022).
[…]
Organization of the Paper.
We start in Section 2 with a precise description of the (KBO) problem, application examples, and implicit differentiation in an RKHS. In Section 3, we describe the empirical (KBO) and state our first main result on the gradient of its value function. Our uniform generalization bound for (KBO) is described in Section 4, together with corollaries for gradient descent and projected gradient descent in Section 4.2. Finally, in Section 5, we discuss the strategy of the proof for the main result.
@@ 2.1 Problem Formulation
We consider the kernel bilevel optimization problem in (KBO) with an RKHS $\mathcal{H}$, which is a space of real-valued functions defined on an input space $\mathcal{X}\subset\mathbb{R}^{p}$ and associated with a reproducing kernel $K:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$. We are interested, in particular, in (regularized) objectives expressed as expectations of point-wise loss functions, a formulation widely adopted in machine learning as it allows the loss functions to represent the average performance over some data distribution. Specifically, given two probability distributions $\mathbb{P}$ and $\mathbb{Q}$ supported on $\mathcal{X}\times\mathcal{Y}$ for some target space $\mathcal{Y}\subset\mathbb{R}^{q}$, we consider objectives of the form:
$\begin{split}L_{out}(\omega,h)&=\mathbb{E}_{\mathbb{Q}}\left[\ell_{out}(\omega,h(x),y)\right],\\ L_{in}(\omega,h)&=\mathbb{E}_{\mathbb{P}}\left[\ell_{in}(\omega,h(x),y)\right]+\frac{\lambda}{2}\|h\|_{\mathcal{H}}^{2},\end{split}$
where $\ell_{in}\text{ and }\ell_{out}:\mathbb{R}^{d}\times\mathbb{R}\times\mathbb{R}^{q}\to\mathbb{R}$ represent the inner and outer point-wise loss functions, $\lambda>0$ is the regularization parameter which is fixed through this work, and $\|\cdot\|_{\mathcal{H}}$ denotes the norm in the RKHS $\mathcal{H}$. The regularization term in $L_{in}$ is often used in practice to prevent overfitting, by penalizing overly complex models. In our setting, it ensures strong convexity of $h\mapsto L_{in}(\omega,h)$, under mild assumptions on $\ell_{in}$, which will be critical to leverage functional implicit differentiation.
Assumptions. Through the paper, we will make the following four assumptions to derive generalization bounds while retaining a simple and modular presentation.
@@
    (Boundedness of $K$). There exists a positive constant $\kappa>0$ such that $K(x,x)\leq\kappa$, for any $x\in\mathcal{X}$.
@@
  4. (D)
    (Convexity of $\ell_{in}$ in its second argument). For any $(\omega,y)\in\mathbb{R}^{d}\times\mathbb{R}^{q}$, the map $v\mapsto\ell_{in}(\omega,v,y)$ is convex.
Assumption (A) on the kernel $K$ holds for a wide class of kernels, such as the Gaussian and Matérn kernels (Wendland, 2004), both of which fulfill this assumption with $\kappa=1$. Assumption (B) on the set $\mathcal{Y}$ is a mild assumption that holds in many practical cases, such as classification, where the set $\mathcal{Y}$ is finite, or when $\mathcal{Y}=[0,1]^{q}$, where complex data, such as images, can be represented. Assumption (C) on the point-wise objectives is a mild regularity assumption. Assumptions (A), (B) and (C) can be relaxed at the expense of weaker yet more technical assumptions, such as finite moments assumptions on $\mathbb{P}$ and $\mathbb{Q}$, and suitable polynomial growth of the kernel and some partial derivatives of $\ell_{in}$ and $\ell_{out}$. It is also sufficient to require that Assumption (C) holds on $\mathcal{U}\times\mathbb{R}\times\mathbb{R}^{q}$ where $\mathcal{U}$ is an open neighborhood of $\mathcal{C}$ and that Assumption (D) holds for any $\omega\in\mathcal{C}$ and $y\in\mathcal{Y}$. We prefer to keep these stronger yet simpler assumptions for clarity. Finally, Assumption (D) is essential to ensure the existence and uniqueness of a smooth minimizer $h^{\star}_{\omega}$. It is a relatively weak assumption that was recently considered in (Petrulionytė et al., 2024) in the context of functional bilevel optimization and that holds in many cases of interest, as discussed in Section 2.2.
2.2 Examples of KBO in Machine Learning
@@
Hyper-Parameter Selection under Distribution Shift. In this application, the aim is to select the best hyper-parameters for a machine learning model, e.g., regularization parameters, while accounting for distribution shift between the training and testing data, i.e., when the training and test data distributions are different (Pedregosa, 2016; Franceschi et al., 2018). This can be viewed as an instance of (KBO) when using models in an RKHS. At the inner-level, the model is trained to minimize the regularized training squared error loss, where the hyper-parameter $\omega>0$ denotes a weight for the data fitting term. At the outer-level, the task is to select the hyper-parameter $\omega$ that minimizes the model’s performance on the distribution-shifted test. Both inner and outer objectives can thus be formulated as:
$L_{out}(\omega,h^{\star}_{\omega})$ $=\mathbb{E}_{x,y}\left[{\left\lvert h^{\star}_{\omega}(x)-y\right\rvert}^{2}\right],$
$L_{in}(\omega,h)$ $=\omega\mathbb{E}_{x,y}\left[{\left\lvert h(x)-y\right\rvert}^{2}\right]+\frac{1}{2}{\left\lVert h\right\rVert}_{\mathcal{H}}^{2}.$
This formulation could be used for domain adaptation (Ben-David et al., 2006) or domain generalization (Wang et al., 2022) to choose hyper-parameters that perform well on the distribution-shifted test data.
Instrumental Variable Regression is a technique used to address endogeneity in statistical modeling by leveraging instruments to estimate causal relationships (Newey & Powell, 2003). The goal here is to estimate a function $t\mapsto f_{\omega}(t)$ parameterized by a vector $\omega$, that satisfies $y=f_{\omega}(t)+\epsilon$, where $y\in\mathbb{R}$ is the observed outcome, $t$ is a treatment, and $\epsilon$ is the error term. The key issue is that $t$ is endogenous, which means that it is correlated with $\epsilon$, i.e., $\mathbb{E}[\epsilon\mid t]\neq 0$, making direct regression inconsistent. Indeed, such correlation leads to biased estimates of $f_{\omega}(t)$ because the assumption of exogeneity, i.e., independence of $t$ and $\epsilon$, is violated. To resolve this, an instrumental variable $x$ that is uncorrelated with $\epsilon$, i.e., $\mathbb{E}[\epsilon\mid x]=0$, but correlated with $t$, can be used to recover the relationship between $y$ and $t$, without being directly affected by the bias introduced by $\epsilon$ using the two stages least squares regression (Singh et al., 2019; Meunier et al., 2024). As shown in (Petrulionytė et al., 2024), this approach can be naturally expressed as a bilevel optimization problem of the form:
$L_{out}(\omega,h^{\star}_{\omega})$ $=\mathbb{E}_{x,y}\left[{\left\lvert h^{\star}_{\omega}(x)-y\right\rvert}^{2}\right],$
$L_{in}(\omega,h)$ $=\mathbb{E}_{x,t}\left[{\left\lvert h(x)-f_{\omega}(t)\right\rvert}^{2}\right]+\frac{\lambda}{2}{\left\lVert h\right\rVert}_{\mathcal{H}}^{2},$

### v3

@@ Abstract
Bilevel optimization has emerged as a technique for addressing a wide range of machine learning problems that involve an outer objective implicitly determined by the minimizer of an inner problem. While prior works have primarily focused on the parametric setting, a learning-theoretic foundation for bilevel optimization in the nonparametric case remains relatively unexplored. In this paper, we take a first step toward bridging this gap by studying Kernel Bilevel Optimization (KBO), where the inner objective is optimized over a reproducing kernel Hilbert space. This setting enables rich function approximation while providing a foundation for rigorous theoretical analysis. In this context, we derive novel finite-sample generalization bounds for KBO, leveraging tools from empirical process theory. These bounds further allow us to assess the statistical accuracy of gradient-based methods applied to the empirical discretization of KBO. We numerically illustrate our theoretical findings on a synthetic instrumental variable regression task.
@@ 1 Introduction
Bilevel optimization involves a nested structure where one optimization problem, called outer-level, is constrained by the solution of another one, called inner-level [candler1977multi]. This formulation has found applications in a broad spectrum of machine learning fields, including hyperparameter tuning [larsen1996design, bengio2000gradient, franceschi2018bilevel], meta-learning [bertinetto2018metalearning, pham2021contextual], inverse problems [holler2018bilevel], and reinforcement learning [hong2023two, liu2023value], making it a powerful tool in theoretical and practical contexts. Its widespread use naturally raises fundamental questions about the generalization properties of models learned through this procedure as the number of data samples increases. Several existing works have studied the generalization and convergence of bilevel algorithms under the assumption that the inner-level problem is strongly convex and that its parameters lie in a finite-dimensional space. These include analyses of the convergence of stochastic bilevel optimization algorithms [Arbel:2021a, Dagreou2022SABA, ghadimi2018approximation, Ji:2021a] and approaches based on algorithmic stability [bao2021stability, zhang2024fine]. The strong convexity assumption ensures a unique inner-level solution, which is crucial for stability and convergence analysis in bilevel optimization. Moreover, restricting the inner-level parameters to a finite-dimensional space does not cover possibly richer infinite-dimensional spaces, as in kernel methods, where the parameter’s dimension may grow with the sample size. This sample size dependence in nonparametric methods poses additional challenges as solutions at different sample sizes are not directly comparable. In contrast, in the finite-dimensional setting, generalization bounds can be derived by quantifying the convergence of inner-level finite-sample solutions toward the infinite samples population solution limit, within a fixed Euclidean space.
@@
Albeit theoretically convenient, strong convexity and finite-dimensionality limit models expressiveness, restricting them to linear functions. Moving beyond linear models requires either relaxing the strong convexity assumption to accommodate more expressive models, such as deep neural networks [Goodfellow-et-al-2016], or considering nonparametric bilevel problems, where the inner-level variable lies in an expressive infinite-dimensional function space, such as a Reproducing Kernel Hilbert Space (RKHS) [scholkopf2002learning]. Early works in bilevel optimization for machine learning followed the latter approach, developing methods for hyperparameter selection in kernel-based models [keerthi2006efficient, kunapuli2008classification]. These works leverage the representer theorem [scholkopf2001generalized] to transform the infinite-dimensional problem into a finite-dimensional one, with dimension depending on the sample size. However, they do not address how sample size impacts generalization. Another line of research instead focuses on relaxing the strong convexity assumption, proposing bilevel algorithms in the absence of convexity [arbel2022nonconvex, kwon2024on, shen2023penalty]. Yet, non-convex bilevel optimization is a very hard problem in general [liu2021towards, arbel2022nonconvex, bolte2024geometric], and strong generalization guarantees in this setting remain out of reach due to instabilities of the inner-level solutions. Overall, learning theory for bilevel problems beyond the strongly convex parametric setting is essentially lacking.
@@
In the present work, we take an initial step toward developing a learning theory that goes beyond the finite-dimensional setting. Specifically, we propose to study Kernel Bilevel Optimization (KBO) problems, where the inner objective $L_{in}:\mathbb{R}^{d}\times\mathcal{H}\rightarrow\mathbb{R}$ finds an optimal inner solution $h_{\omega}^{\star}$ in an RKHS $\mathcal{H}$ for a given parameter $\omega$ in $\mathbb{R}^{d}$, while the outer objective $L_{out}:\mathbb{R}^{d}\times\mathcal{H}\rightarrow\mathbb{R}$ optimizes the parameter $\omega$ over a closed subset $\mathcal{C}$ of $\mathbb{R}^{d}$, given the inner solution $h_{\omega}^{\star}$:
$\min_{\omega\in\mathcal{C}}\mathcal{F}(\omega)\coloneqq L_{out}(\omega,h^{\star}_{\omega})\quad\text{s.t.}\quad h^{\star}_{\omega}=\operatorname*{arg\,min}_{h\in\mathcal{H}}L_{in}(\omega,h).$ (KBO)
@@
In particular, we focus on objectives that are expectations of point-wise losses, a common setting in learning theory. RKHS provides a natural framework to study learning-theoretic arguments, and has been instrumental for many fruitful results in pattern recognition and machine learning. They allow to describe very expressive non-linear models with simple and stable algorithms, while enabling a rich statistical analysis and featuring adaptivity to the regularity of the population problem [shawe2004kernel, scholkopf2002learning, hofmann2008kernel]. Our choice is also motivated by the relevance of kernel methods, even in the deep learning era. They remain competitive for some prediction problems, such as those involving physics [doumeche2024physics, letizia2022learning]. Additionally, the mathematics of kernel methods are useful to describe the limiting behavior of deep network training for very large models [jacot2018neural, belkin2018understand]. In this limit, the problem becomes (strongly) convex in an infinite-dimensional function space, simplifying the difficulties of non-convex model parameterizations, a major bottleneck in the analysis of such models. This point of view was leveraged by petrulionyte2024functional who introduced functional bilevel optimization, and our setting can be seen as a special case for which the underlying function space is an RKHS. From a practical perspective, our setting is amenable to first-order methods using implicit differentiation techniques [griewank2003piggyback, bai2019deep, blondel2022efficient].
Contributions. We leverage empirical process theory and its extension to $U$-processes [sherman1994maximal] to derive uniform generalization bounds for the value function of (KBO), quantifying the discrepancy between $\mathcal{F}$ and its plug-in estimator $\widehat{\mathcal{F}}$ in terms of both their values and gradients. Classical empirical process results [vaart_wellner_1996] do not directly apply here, as our functional setting involves processes taking values in an infinite-dimensional space rather than real numbers. To address this, we exploit the RKHS structure to represent the resulting error terms as real-valued $U$-processes, enabling the use of results from sherman1994maximal. The control in terms of gradients is crucial to study first-order optimization methods, since $\mathcal$$\widehat${F}} is typically non-convex and iterative methods seek to find approximate critical points where $\|\nabla\widehat{\mathcal{F}}\|$ is small. Our result relies on an equivalence we establish between $\nabla\widehat{\mathcal{F}}$ and a plug-in statistical estimate of $\nabla\mathcal{F}$ that is more amenable to a statistical analysis. We then use our uniform bounds to provide generalization guarantees for gradient descent and projected gradient descent applied to $\widehat{\mathcal{F}}$. Under specific assumptions, we show convergence rates for sub-optimality measures that depend on the sample sizes and the number of algorithmic iterations. This illustrates the practical relevance of our generalization bounds on simple bilevel algorithms. For a large number of steps, gradient algorithms applied to the empirical (KBO) find approximate critical points of the population (KBO) up to a statistical error which we control.
Organization of the paper. In Section˜2, we describe (KBO), give two application examples, and explain implicit differentiation in an RKHS. In Section˜3, we present the empirical (KBO) and state our first main result on the gradient of its value function. Section˜4 provides uniform generalization bounds for (KBO), with applications to bilevel gradient methods, as well as a sketch of the proof of our main result. Finally, in Section˜5, we illustrate our theoretical findings with experiments on synthetic data for the instrumental variable regression problem.
@@ 2.1 Problem formulation
We consider the (KBO) problem with an RKHS $\mathcal{H}$, which is a space of real-valued functions defined on a Borel input space $\mathcal{X}\subset\mathbb{R}^{p}$ and associated with a reproducing kernel $K:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$. We are interested, in particular, in (regularized) objectives expressed as expectations of point-wise loss functions, a formulation widely adopted in machine learning as it allows the loss functions to represent the average performance over some data distribution. Specifically, given two probability distributions $\mathbb{P}$ and $\mathbb{Q}$ supported on $\mathcal{X}\times\mathcal{Y}$ for some target space $\mathcal{Y}\subset\mathbb{R}^{q}$, we consider objectives of the form:
$L_{out}(\omega,h)=\mathbb{E}_{\mathbb{Q}}\left[\ell_{out}(\omega,h(x),y)\right],\quad L_{in}(\omega,h)=\mathbb{E}_{\mathbb{P}}\left[\ell_{in}(\omega,h(x),y)\right]+\frac{\lambda}{2}\|h\|_{\mathcal{H}}^{2},$
where $\ell_{in}\text{ and }\ell_{out}:\mathbb{R}^{d}\times\mathbb{R}\times\mathbb{R}^{q}\to\mathbb{R}$ represent the inner and outer point-wise loss functions, $\lambda>0$ is the regularization parameter which is fixed through this work, and $\|\cdot\|_{\mathcal{H}}$ denotes the norm in the RKHS $\mathcal{H}$. The regularization term in $L_{in}$ is often used in practice to prevent overfitting by penalizing overly complex models. In our setting, it ensures strong convexity of $h\mapsto L_{in}(\omega,h)$ under mild assumptions on $\ell_{in}$, which will be critical to leverage functional implicit differentiation.
Assumptions. Through the paper, we make the following five assumptions to derive generalization bounds while retaining a simple and modular presentation.
@@
    (Measurability of $K$). $K$ is measurable on $\mathcal{X}\times\mathcal{X}$.
@@
    (Boundedness of $K$). There exists a constant $\kappa>0$ such that $K(x,x)\leq\kappa$, for any $x\in\mathcal{X}$.
  3. (C)
@@
  5. (E)
    (Convexity of $\ell_{in}$). For any $(\omega,y)\in\mathbb{R}^{d}\times\mathbb{R}^{q}$, the map $v\mapsto\ell_{in}(\omega,v,y)$ is convex.
Assumptions˜(A) and (B) on $K$ hold for a wide class of kernels, such as the Gaussian, Laplacian, and Matérn [rasmussen2006gaussian] kernels. They also hold if $K$ is a Mercer kernel [mercer1909functions], i.e., a continuous, positive‐definite kernel on a compact domain $\mathcal{X}$. Specifically, a kernel built using neural network features, e.g., neural tangent kernel [jacot2018neural], satisfies these assumptions on the space of images, which is compact since the pixel values have a bounded range. Assumption˜(C) on $\mathcal{Y}$ is a mild assumption that holds in most supervised learning applications, such as classification where $\mathcal{Y}$ is finite, or cases where $\mathcal{Y}=[0,1]^{q}$, enabling the representation of complex data, like images. Assumption˜(D) on the point-wise objectives is a mild regularity assumption, which is met for the most commonly used loss functions in practice, including the squared loss, logistic loss, cross-entropy loss, and KL divergence. Finally, Assumption˜(E) is essential to ensure the existence and uniqueness of a smooth minimizer $h^{\star}_{\omega}$. It is a relatively weak assumption that was recently considered in [petrulionyte2024functional] in the context of functional bilevel optimization, and that holds in many cases of interest, as discussed in Section˜2.2.
Remark 2.1.
Assumptions˜(B), (C) and (D) can be relaxed at the expense of weaker yet more technical assumptions, such as finite moment assumptions on $\mathbb{P}$ and $\mathbb{Q}$, and suitable polynomial growth of the kernel and some partial derivatives of $\ell_{in}$ and $\ell_{out}$. It is also sufficient to require that Assumption˜(D) holds on $\mathcal{U}\times\mathbb{R}\times\mathbb{R}^{q}$ where $\mathcal{U}$ is an open neighborhood of $\mathcal{C}$, and that Assumption˜(E) holds for any $\omega\in\mathcal{C}$ and $y\in\mathcal{Y}$. We prefer to keep these stronger yet simpler assumptions for clarity.
2.2 Examples of (KBO) in machine learning
@@
Hyperparameter selection under distribution shift. In this application, the aim is to select the best hyperparameters for a machine learning model, e.g., regularization parameters, while accounting for distribution shift between the training and test data, i.e., when the training and test data distributions $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$ are different [pedregosa2016hyperparameter, franceschi2018bilevel]. This can be viewed as an instance of (KBO) when using models in an RKHS. At the inner-level, the model $h$ is trained to minimize the regularized training squared error loss, with the hyperparameter $\omega>0$ representing the weight for the data fitting term. At the outer-level, the task is to select the hyperparameter $\omega$ that maximizes the model’s performance on the distribution-shifted test data. Both inner and outer objectives can thus be formulated as:
$L_{out}(\omega,h)=\frac{1}{2}\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{test}}}\left[\left\lvert h(x)-y\right\rvert^{2}\right],\quad L_{in}(\omega,h)=\frac{\omega}{2}\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{train}}}\left[\left\lvert h(x)-y\right\rvert^{2}\right]+\frac{\lambda}{2}\left\lVert h\right\rVert_{\mathcal{H}}^{2}.$
This formulation could be used for domain adaptation [ben2006analysis] or domain generalization [wang2022generalizing] to choose hyperparameters that perform well on the distribution-shifted test data.
Instrumental variable regression. It is a technique used to address endogeneity in statistical modeling by leveraging instruments to estimate causal relationships [newey2003instrumental]. The goal is to estimate a function $t\mapsto f_{\omega}(t)$ parameterized by a vector $\omega$, that satisfies $y=f_{\omega}(t)+\epsilon$, where $y\in\mathbb{R}$ is the observed outcome, $t$ is the treatment, and $\epsilon$ is the error term. The key issue is that $t$ is endogenous, which means that it is correlated with $\epsilon$, making direct regression inconsistent. Indeed, such correlation leads to biased estimates of $f_{\omega}(t)$ as the assumption of exogeneity, i.e., independence of $t$ and $\epsilon$, is violated. To resolve this, one can use an instrumental variable $x$, uncorrelated with $\epsilon$ but correlated with $t$, to recover the relationship between $y$ and $t$ without being directly affected by the bias introduced by $\epsilon$, typically via two-stage least squares regression [singh2019kernel, meunier2024nonparametric]. As shown in [petrulionyte2024functional], this approach can be naturally expressed as a bilevel problem with inner and outer objectives of the form:
$L_{out}(\omega,h)=\frac{1}{2}\mathbb{E}_{x,y}\left[\left\lvert h(x)-y\right\rvert^{2}\right],\quad L_{in}(\omega,h)=\frac{1}{2}\mathbb{E}_{x,t}\left[\left\lvert h(x)-f_{\omega}(t)\right\rvert^{2}\right]+\frac{\lambda}{2}\left\lVert h\right\rVert_{\mathcal{H}}^{2},$